# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v2

@@ 2.2 Risk Minimization Principles
Such a principle 35 effectively improves robustness in distribution shifts and has shown positive effects on foundation models 36, 37. It constructs a collection of uncertainty sets over tasks and results in the optimization objective as follows:
@@ 3 Results
Previous robust adaptation works require model evaluation over task batches for robust optimization while the task batch size $\mathcal{B}$ is restricted by the sample or memory cost. Meanwhile, nearly all of them fail to utilize the above optimization outcomes $H_{t}$; our finding is that under Assumptions 1/2/3, the cumulated risk episodes can be used to train a risk predictive model, i.e., the risk learner in MPTS, through posterior inference and work as the informative prior for continually scoring task difficulties.
@@
MPTS amortizes arbitrary task adaptation performance evaluation and approximately scores task difficulties. Learning to predict the optimization outcome $p(\ell|\bm{\tau},H_{1:t};\bm{\theta}_{t})$ is hopeful for performance evaluation in infinite tasks with negligible computations and enlarges the pseudo batch size $\hat{\mathcal{B}}$ for filtering, implicitly encouraging exploration in the task space. Theorem 1 provides a provable foundation for optimization outcome prediction, which suggests that a controllable perturbation over $\bm{\theta}$, such as one-step gradient update, changes the relative difficulties between tasks with minor chance. Hence, the candidate tasks $\mathcal{T}_{t+1}^{\hat{\mathcal{B}}}$ evaluated by the latest machine learner $\bm{\theta}_{t}$ in posterior inference preserves their relative task difficulties in the case given its update $\bm{\theta}_{t+1}$. The following empirical results will confirm such predictability and its difficulty scoring ability to some extent with MPTS.
@@
The experimental design takes the benchmark typicality and the practical challenges into consideration. Testing scenarios span downstream tasks across pattern recognition and decision-making, with certain experiments incorporating multimodal foundation models such as CLIP backbones to assess MPTS. These experiments mainly examine few-shot adaptation and include (1) K-shot sinusoid regression 28, (2) N-way K-shot image classification 42 with CLIP models and (3) Meta-RL 28. Additionally, MPTS explores the application scenarios like (4) robotic DR 43, 9 for zero-shot adaptation and (5) CLIP model’s parameter-efficient finetuning (PEFT) towards image classification.
@@
Note that the primary focus of this study lies in boosting adaptation robustness; the comparison baselines are in accordance with the previously mentioned risk minimization principles. While these principles, along with our MPTS, are agnostic to specific zero-shot, few-shot, or many-shot learning methods, we adopt the representative or state-of-the-art (SOTA) method as the default backbone in the experiment.
In sinusoid regression and Meta-RL [^1] To ensure enough coverage of task space, we adopt a mixture strategy of MPTS and random sampling as an empirical regularizer in all RL scenarios. Please refer to Section A for details., we employ MAML 28 as the backbone algorithm. In N-way K-shot image classification, CLIP 1 has already achieved remarkable superior zero-shot performance, and we utilize its prompt-based extension MaPLe 44 for classification. In robotic DR for zero-shot adaptation, we set TD3 45 as the base policy optimization algorithm due to its stable performance. In many-shot adaptation, we still use MaPLe 44 for prompt-tuning in image classification.
To implement baselines, we incorporate Table 1 listed ERM 34, DRM 16, 33, GDRM 35 and MPTS (ours) into the above backbone methods. In terms of adaptation robustness, we compute $\text{CVaR}_{\alpha}$ across validation task set or testing task datasets with $\alpha=\{0.9,0.7,0.5\}$, with some OOD results reported. Moreover, we assess these methods from efficiency metrics, such as computational cost, memory usage, and task efficiency in separate scenarios. For a fair comparison, the task batch to optimize is the same for all baselines, excluding those easier tasks pruned after evaluation. Unless notable clarifications, the default task batch size for ERM and GDRM is $\mathcal{B}$ while that for DRM is ^$\hat{\mathcal{B}}=2\mathcal{B}$ with half filtered for stable optimization. We refer the reader to Supplementary Notes Section F/G for all configuration details.
@@ 3.1 Demonstration of the MPTS’s role in K-shot sinusoid regression
In challenging tasks’ prioritized sampling 14, 15, 16, assessing task difficulties is identified as the preliminary step. Note that the risk learner learns to amortize the exact evaluation $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}_{t})$ $\forall\bm{\tau}\in\mathcal{T}\ \text{and}\ \bm{\theta}_{t}\in\bm{\Theta}$ with the help of risk histories and the evaluation result implicitly reflects adaptation difficulties. Once the risk learner can roughly compare difficulties between tasks, we can trust the amortized result in MPTS for the follow-up robust optimization.
Hence, given $(t+1)$-th screened task batch and its evaluation, we keep track of the risk learner’s predicted adaptation risk values, {¯$\{\bar{\ell}_{t+1,i}:\approx\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}[p_{\bm{\psi}}(\ell|\bm{\tau}_{t+1,i},H_{1:t})]\}_{i=1}^{\mathcal{B}}$ and the corresponding exact ones $\{\ell_{t+1,i}\}_{i=1}^{\mathcal{B}}$ from $\bm{\theta}_{t+1}$ for computing the Pearson correlation coefficient (PCC) value $\rho_{\bar{\ell},\ell}:=\frac{\sum_{i=1}^{\mathcal{B}}(\bar{\ell}_{t+1,i}-\text{Mean}[\{\bar{\ell}_{t+1,.}\}])(\ell_{t+1,i}-\text{Mean}[\{\ell_{t+1,.}\}])}{\sqrt{\sum_{i=1}^{\mathcal{B}}(\bar{\ell}_{t+1,i}-\text{Mean}[\{\bar{\ell}_{t+1,.}\}])^{2}}\sqrt{\sum_{i=1}^{\mathcal{B}}(\ell_{t+1,i}-\text{Mean}[\{\ell_{t+1,.}\}])^{2}}}$ over iterations. The predicted risk values are actually amortized evaluation results conditioned on $\bm{\theta}_{t}$, but it will not affect empirical conclusion from PCC values due to the provably relative difficulty relationship preservation in Theorem 1. Under this consideration, PCC values reasonably quantify the difficulty-scoring capability of the risk learner in a rough granularity.
@@
The risk learner accelerates the learning process and improves comphrehensive adaptation performance with the help of pseudo-evaluation.
@@
In meta-testing, Fig. 3b confirms the MPTS and DRM’s superior average and $\text{CVaR}_{\alpha}$ performance, and their advantages over GDRM and ERM grow significant with increasing confidence level $\alpha$. As empirically verified in work 46, DRM typically sacrifices task and computational efficiency to enhance the robustness of machine learners, often requiring intensive task evaluation to filter. For the used MAML backbone, the gradient update as the inner loop for adaptation further increases the computational overhead, while MPTS implicitly bypasses it through prediction.
@@
We first examine the acquisition function $\mathcal{A}(\mathcal{T}^{\mathcal{B}};\bm{\phi},\bm{\psi})$’s influence through the trade-off parameters $\{\gamma_{0},\gamma_{1}\}$ in Eq. (16). To this end, we meta-test machine learners’ average performance trained with various setups $\frac{\gamma_{1}}{\gamma_{0}}=\{1.0,3.0,6.0,9.0\}$, $\gamma_{0}=0.0$ and $\gamma_{1}=0.0$ in Fig. 3c. The results show that relatively higher uncertainty weights can decrease the average MSEs to a certain level while simply dropping out the worst-case term with $\gamma_{0}=0.0$ weakens the performance.
Moreover, we scrutinize the influence of the pseudo batch size $\hat{\mathcal{B}}$ in Fig. 3d, where testing models are trained with diverse $\hat{\mathcal{B}}$. Increasing $\hat{\mathcal{B}}$ to a certain scale significantly reduces the average meta-testing MSEs in Fig. 3d. Nevertheless, excessively larger $\hat{\mathcal{B}}$ configuration, e.g., $8\times\mathcal{B}$, adversely affects testing performance. This is because the selected task identifier tends to concentrate under the worst-case selection principle when increasing the number of identifiers and over-optimizing the local region of the task space deteriorates the global generalization.
Overall, the appropriate configuration of MPTS adheres to the following disciplines: (i) $\hat{\mathcal{B}}$ greater than $\mathcal{B}$ should be set within an acceptable size to encourage task exploration and prevent excessive optimization within a local task region; (ii) since adaptation robustness improve remains the primary goal, we consistently set $\gamma_{0}\in\mathbb{R}^{+}$ as the default across all experiments.
Figure 4: 5-way 1-shot Meta-testing Classification Results. a-d. Shown are testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with meta-trained machine learners on different datasets. e. With experiments on ImageNet-A as an example, we report the memory cost and clock time relative to ERM during meta-training for all methods.
@@ 3.2 Few-Shot adaptation benefits from MPTS in robustness and learning efficiency
Prompt-based N-way K-shot image classification.
@@
We perform 5-way 1-shot image classification performance using MaPLe. The six meta-training datasets are respectively constructed by ImageNet-CG 47, ImageNet-CI 47, ImageNet-CS 47, ImageNet-A 48, ImageNet-S 49 and ImageNet-R 50 in a standard manner. Explicit $\bm{\tau}$ are unavailable to specify the task; however, this can be approximately overcome by representing the identifier through a small reference model. Specifically, we leverage CLIP’s text encoder to obtain $\bm{\tau}\approx[\text{CLIP}_{\text{text}}(\mathcal{C}_{1}),\dots,\text{CLIP}_{\text{text}}(\mathcal{C}_{K})]$ with the tokenizations of $K$ class texts $\mathcal{C}_{1:K}$ from $\mathcal{D}_{\tau}^{S}$.
@@
During meta-training, resource consumption on large image datasets is quantified in Fig. 4e, comparing relative computational time and memory usages across methods. The extra resource cost arising from optimizing risk learners can be negligible in MPTS compared with the total cost. DRM demands relatively 1.3x computational time and 1.6x memory consumption compared with ERM, as $2\mathcal{B}$ tasks are fed into the model to evaluate and filter half in each iteration.
@@
In meta-testing, MPTS achieves the highest average accuracies on six image datasets in Fig. 4d and Table 5. As for robustness, the results in Fig. 4a-c witnesses that both MPTS and DRM can comprehensively increase $\text{CVaR}_{0.5}$, $\text{CVaR}_{0.7}$ and $\text{CVaR}_{0.9}$ values. For a specific method, its performance exhibits similar patterns across different datasets. All robust optimization methods show robustness improvement over ERM to some extent. Among them, MPTS and DRM are leading in all metrics, while the latter requires more learning resources to guarantee adaptation robustness.
@@
Figure 5: Meta-RL Results on Five Mujoco Environments (7 Runs). a. The cumulative returns with standard error of means (SEMs) belonging to $\text{CVaR}_{0.9}$ validation MDPs are displayed during meta-training. b. We compute the average cumulative returns with SEMs on all validation MDPs during meta-training. c. Tracked are the risk learner’s PCC values with SEMs over training iterations. d. The relative clock time quantifies the computational complexity for all methods on Walker2dVel as an example, where the computational time of ERM works as the anchor. e. We report returns at $\alpha$ levels of $\text{CVaR}_{\alpha}$ on meta-testing MDPs. f. The box-plot reports meta-testing results averaged over meta-testing MDPs.
Meta reinforcement learning.
@@
We first analyze the meta-training results. As displayed in Fig. 2a, the identifier refers to specific physics variables and is explicit for risk learners’ construction. In Fig. 5a, on all benchmarks except ReacherPos, MPTS reaches the highest $\text{CVaR}_{0.9}$ validation returns, followed by DRM and GDRM. ERM largely underperforms MPTS, which is consistent on all benchmarks. In terms of computational and sample efficiency, DRM takes about 1.5x computational time on Walker2dVel due to extra policy evaluation to filter tasks after meta gradient updates, while MPTS does not suffer this issue (see Fig. 5d). In Fig. 5b, DRM sacrifices more average returns than ERM on HalfCheetahMassVel, HalfCheetahVel, and Walker2dVel. Though MPTS can be viewed as the surrogate for DRM, the involvement of sampling randomness and task exploration benefit ensure MPTS’s average performance is comparable with ERM on HalfCheetahMassVel and Walker2dVel. GDRM behaves mediate throughout evaluations in Fig. 5a-b. DRM mostly strikes a balance between the average and $\text{CVaR}_{0.9}$ performance; however, DRM and MPTS receive both the average and $\text{CVaR}_{0.9}$ bonus on ReacherPos. This reflects that the effect of worst task selection might vary between environments. The tracked $\rho_{\bar{\ell},\ell}$ in Fig. 5c witness the risk learner’s strong capability of discriminating MDP difficulties, staying above 0.5 nearly across all environments over iterations. The trend of $\rho_{\bar{\ell},\ell}$ on ReacherPos differs from those on other environments and does not decline along iteration.
@@
Meta-testing results in Fig. 5e-f reveals performance differences across CVaR values and robustness improvements of other baselines over ERM, which align with the learning curves in Fig. 5a. MPTS’s performance advantage notably amplifies with increasing $\alpha$ values. In the extreme worst case $\text{CVaR}_{0.9}$, MPTS beats ERM by more than 20% returns on all benchmarks. Average meta-testing performance across environments draws no consistent conclusions. Walker2dMassVel and Walker2dVel show minor differences between methods (see Fig. 5f), while on HalfCheetahMassVel, MPTS achieves comparable returns to ERM despite slightly higher variance. HalfCheetah verifies the marginal advantages of GDRM and ERM, whereas ReacherPos favors MPTS and DRM with reduced variance. Importantly, MPTS, serving as a DRM proxy, essentially maintains average performance comparable with ERM.
@@
In total, MPTS enjoys significant adaptation robustness benefits across most scenarios with mediate average performance, but it reserves more computational and task efficiency than DRM. Unlike few-shot classification cases, Meta-RL mostly trades off the worst-case and average performance, similarly observed in work 33.
@@
3.3 MPTS still achieves adaptation robustness in zero-shot continuous control

### v3

@@ 2.2 Risk Minimization Principles
Such a principle 35 effectively improves robustness in distribution shifts and has shown positive effects on training foundation models 36, 37. It constructs a collection of uncertainty sets over tasks and results in the optimization objective as follows:
@@ 3 Results
Prior robust adaptation works 14, 15, 16, 17, 20 require evaluating the machine learner over tasks for further robust optimization while the task batch size $\mathcal{B}$ is restricted by the sample or memory cost. Meanwhile, nearly all of them fail to utilize the above optimization outcomes $H_{1:t}$; our finding is that under Assumptions 1/2/3, the cumulated risk episodes can be used to train a risk predictive model, i.e., the risk learner in MPTS, through posterior inference and work as the informative prior for continually scoring each task’s adaptation difficulty.
@@
MPTS amortizes arbitrary task adaptation performance evaluation and approximately ranks task difficulties. Learning to predict the optimization outcome $p(\ell|\bm{\tau},H_{1:t};\bm{\theta}_{t})$ is hopeful for performance evaluation in infinite tasks with negligible computations and enlarges the pseudo batch size $\hat{\mathcal{B}}$ for filtering, implicitly encouraging exploration in the task space. Theorem 1 provides a provable foundation for optimization outcome prediction, which suggests that a controllable perturbation over $\bm{\theta}$, such as a one-step gradient update with a smaller enough learning rate, changes the overall difficulties between tasks with a minor probability. Hence, the candidate tasks $\mathcal{T}_{t+1}^{\hat{\mathcal{B}}}$ evaluated by the latest machine learner $\bm{\theta}_{t}$ in posterior inference probably preserves their relative task difficulties given the updated $\bm{\theta}_{t+1}$. The following empirical results will confirm such predictability and its difficulty scoring ability with MPTS.
@@
The experimental design takes the benchmark typicality and the practical challenges into consideration. Testing scenarios span downstream tasks across pattern recognition and decision-making to assess MPTS, with certain experiments involving multimodal foundation models. These experiments mainly examine few-shot adaptation and include (1) K-shot sinusoid regression 28, (2) N-way K-shot image classification 42 with CLIP models and (3) Meta-RL 28. Additionally, MPTS explores the application scenarios like (4) robotic DR 43, 9 for zero-shot adaptation and (5) CLIP model’s parameter-efficient finetuning (PEFT) towards image classification.
@@
Note that the primary focus of this study lies in boosting adaptation robustness; the comparison baselines are in accordance with the previously mentioned risk minimization principles. While these principles, along with our MPTS, are agnostic to specific zero-shot, few-shot, or many-shot learning methods, we adopt the representative or SOTA method as the default backbone in the experiment.
In sinusoid regression and Meta-RL, we employ MAML 28 as the backbone algorithm. In N-way K-shot image classification, CLIP 1 has already achieved remarkable zero-shot performance, and we utilize its prompt-based extension MaPLe 44 for classification. In robotic DR for zero-shot adaptation, we set TD3 45 as the base policy optimization algorithm due to its stable performance. In many-shot adaptation, we still use MaPLe 44 for prompt-tuning in image classification.
To implement baselines, we incorporate Table 1 listed ERM 34, DRM 16, 33, GDRM 35 and MPTS (ours) into the above backbone methods. In terms of adaptation robustness, we compute $\text{CVaR}_{\alpha}$ across validation task set or testing task datasets with $\alpha=\{0.9,0.7,0.5\}$, with some OOD results reported. Moreover, we evaluate these methods from efficiency metrics, such as computational cost, memory usage, and sample efficiency in separate scenarios. For a fair comparison, the task batch to optimize $\mathcal{B}$ is the same for all baselines, excluding those easier tasks pruned after evaluation. Unless notable clarifications, the sampled task batch size for ERM and GDRM is $\mathcal{B}$ while that for DRM is ^$\hat{\mathcal{B}}=2\mathcal{B}$ with half filtered for stable optimization. We refer the reader to Supplementary Notes Section F/G for configuration details.
@@ 3.1 Demonstration of the MPTS’s role in K-shot sinusoid regression
In challenging tasks’ prioritized sampling 14, 15, 16, 33, assessing task difficulties is identified as the preliminary step. Note that the risk learner learns to amortize the exact evaluation $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}_{t})$ $\forall\bm{\tau}\in\mathcal{T}\ \text{and}\ \bm{\theta}_{t}\in\bm{\Theta}$ with the help of risk histories, and the evaluation result implicitly reflects adaptation difficulties. Once the risk learner can roughly judge difficulties between tasks, we can trust the amortized result in MPTS for the follow-up robust optimization.
Hence, given $(t+1)$-th screened task batch and its evaluation, we keep track of the risk learner’s predicted adaptation risk values, {¯$\{\bar{\ell}_{t+1,i}:\approx\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}[p_{\bm{\psi}}(\ell|\bm{\tau}_{t+1,i},H_{1:t})]\}_{i=1}^{\mathcal{B}}$ and the corresponding exact ones $\{\ell_{t+1,i}\}_{i=1}^{\mathcal{B}}$ from $\bm{\theta}_{t+1}$ for computing the Pearson correlation coefficient (PCC) value $\rho_{\bar{\ell},\ell}:=\frac{\sum_{i=1}^{\mathcal{B}}(\bar{\ell}_{t+1,i}-\text{Mean}[\{\bar{\ell}_{t+1,.}\}])(\ell_{t+1,i}-\text{Mean}[\{\ell_{t+1,.}\}])}{\sqrt{\sum_{i=1}^{\mathcal{B}}(\bar{\ell}_{t+1,i}-\text{Mean}[\{\bar{\ell}_{t+1,.}\}])^{2}}\sqrt{\sum_{i=1}^{\mathcal{B}}(\ell_{t+1,i}-\text{Mean}[\{\ell_{t+1,.}\}])^{2}}}$ over iterations. The predicted risk values are actually amortized evaluation results conditioned on $\bm{\theta}_{t}$, but it will not affect empirical conclusion from PCC values due to the provably relative difficulty relationship preservation in Theorem 1. Under this consideration, PCC values reasonably quantify the task difficulty ranking capability of the risk learner in a rough granularity.
@@
The risk learner accelerates the learning process and improves comphrehensive adaptation performance with the help of amortized evaluation.
@@
In meta-testing, Fig. 3b confirms the MPTS and DRM’s lowest average and $\text{CVaR}_{\alpha}$ mean-square errors (MSEs), and their advantages over GDRM and ERM grow significant with increasing confidence level $\alpha$. As empirically verified in work 46, DRM typically sacrifices task and computational efficiency to enhance the robustness of machine learners, often requiring intensive task evaluation to filter. For the used MAML backbone, the gradient update as the inner loop for adaptation further increases the computational overhead, while MPTS implicitly bypasses it through probabilistic prediction.
@@
We first examine the acquisition function $\mathcal{A}(\mathcal{T}^{\mathcal{B}};\bm{\phi},\bm{\psi})$’s influence through the trade-off parameters $\{\gamma_{0},\gamma_{1}\}$ in Eq. (16). To this end, we meta-test machine learners’ average performance trained with various setups $\frac{\gamma_{1}}{\gamma_{0}}=\{1.0,3.0,6.0,9.0\}$, $\gamma_{0}=0.0$ and $\gamma_{1}=0.0$ in Fig. 3c. The results depict that relatively higher uncertainty weights can decrease the average MSEs to a certain level while simply eliminating the worst-case consideration with $\gamma_{0}=0.0$ weakens the performance.
Moreover, we scrutinize the influence of the pseudo batch size $\hat{\mathcal{B}}$ in Fig. 3d, where testing models are trained with diverse $\hat{\mathcal{B}}$. Increasing $\hat{\mathcal{B}}$ to a certain scale significantly reduces the average meta-testing MSEs in Fig. 3d. Nevertheless, excessively larger $\hat{\mathcal{B}}$, e.g., ^$\hat{\mathcal{B}}=8\times\mathcal{B}$, adversely affects testing performance. This is because the selected task identifiers might concentrate under the worst-case selection principle when increasing the number of identifiers and over-optimizing the local region of the task space deteriorates the global generalization.
Overall, the appropriate configuration of MPTS adheres to the following disciplines: (i) $\hat{\mathcal{B}}$ greater than $\mathcal{B}$ should be set within an acceptable size to encourage task exploration and prevent excessive optimization within a local task region; (ii) since adaptation robustness improvement remains the primary goal, we consistently set $\gamma_{0}\in\mathbb{R}^{+}$ as the default across all experiments.
Figure 4: 5-way 1-shot Meta-testing Classification Results (3 Runs in Average). a-d. Shown are testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with meta-trained machine learners on different datasets. e. With experiments on ImageNet-A as an example, we report the memory cost and clock time relative to ERM during meta-training for all methods.
@@ 3.2 Few-Shot adaptation benefits from MPTS in robustness and learning efficiency
Result analysis in N-way K-shot image classification.
@@
We perform 5-way 1-shot image classification using MaPLe. The six meta-training datasets are respectively constructed by ImageNet-CG 47, ImageNet-CI 47, ImageNet-CS 47, ImageNet-A 48, ImageNet-S 49 and ImageNet-R 50 in a standard manner. Explicit $\bm{\tau}$ are unavailable to specify the task; however, it can be approximately resolved by describing the identifier through a small reference model. Specifically, we leverage CLIP’s text encoder to obtain $\bm{\tau}\approx[\text{CLIP}_{\text{text}}(\mathcal{C}_{1}),\dots,\text{CLIP}_{\text{text}}(\mathcal{C}_{K})]$ with the tokenizations of $K$ class texts $\mathcal{C}_{1:K}$ from $\mathcal{D}_{\tau}^{S}$.
@@
During meta-training, resource consumption on large image datasets is quantified in Fig. 4e, comparing relative computational time and memory usage across methods. The extra memory and computation cost arising from optimizing risk learners can be negligible in MPTS compared with the total ones. DRM demands relatively 1.3x computational time and 1.6x memory consumption compared with ERM, as $2\mathcal{B}$ tasks are fed into the model to evaluate and filter half in each iteration.
@@
In meta-testing, MPTS achieves the highest average accuracies on six image datasets in Fig. 4d and Table 5. As for robustness, the results in Fig. 4a-c witness a comprehensively increase of $\text{CVaR}_{0.5}$, $\text{CVaR}_{0.7}$ and $\text{CVaR}_{0.9}$ values in both MPTS and DRM. Each specific method’s performance exhibits alike patterns across different datasets. All robust optimization methods show robustness improvement over ERM to some extent. Among them, MPTS and DRM are leading in all metrics, while the latter requires more learning resources to guarantee adaptation robustness. Overall, difficult task prioritization in this benchmark can simultaneously increase average performance as a side product.
@@
Figure 5: Meta-RL Results on Five Mujoco Environments (7 Runs). a. The cumulative returns with standard error of means (SEMs) belonging to $\text{CVaR}_{0.9}$ validation MDPs are displayed during meta-training. b. We compute the average cumulative returns with SEMs on validation MDPs during meta-training. c. Tracked are the risk learner’s PCC values with SEMs over training iterations. d. The relative clock time quantifies the computational complexity for all methods on Walker2dVel, where ERM’s runtime works as the anchor. e. We report $\text{CVaR}_{\alpha}$ returns of meta-testing MDPs. f. The box-plot reports results averaged over meta-testing MDPs.
Result analysis in Meta-RL.
@@
We first analyze the meta-training results. As displayed in Fig. 2a, the identifier refers to specific physics variables, explicit for risk learners’ construction. In Fig. 5a, on all benchmarks except ReacherPos, MPTS reaches the highest $\text{CVaR}_{0.9}$ validation returns, followed by DRM and GDRM. ERM largely underperforms MPTS, which is consistent on all benchmarks. In terms of computational and sample efficiency, DRM takes about 1.5x computational time on Walker2dVel due to extra environment interactions under inner gradient updated policies to filter tasks, while MPTS does not suffer this issue (see Fig. 5d). In Fig. 5b, DRM sacrifices more average returns than ERM on HalfCheetahMassVel, HalfCheetahVel, and Walker2dVel. Though MPTS can be viewed as the surrogate for DRM, the involvement of active sampling and extra task exploration makes MPTS’s average performance comparable with ERM on HalfCheetahMassVel and Walker2dVel. GDRM behaves mediate throughout evaluations in Fig. 5a-b. DRM mostly strikes a balance between the average and $\text{CVaR}_{0.9}$ performance; however, DRM and MPTS receive both the average and $\text{CVaR}_{0.9}$ bonus on ReacherPos. This reflects that the effect of worst task selection might vary between environments. The tracked $\rho_{\bar{\ell},\ell}$ in Fig. 5c reveals the risk learner’s strong capability of discriminating MDP difficulties, staying above 0.5 nearly across all environments over iterations. The trend of $\rho_{\bar{\ell},\ell}$ on ReacherPos differs from those on other environments and does not decline along iteration.
@@
Meta-testing results in Fig. 5e-f displays performance differences across CVaR values and other baselines’ robustness improvements over ERM, which align with the learning curves in Fig. 5a. MPTS’s performance advantage notably amplifies with increasing $\alpha$ values. In the extreme worst case $\text{CVaR}_{0.9}$, MPTS beats ERM by more than 20% returns on all benchmarks. Average meta-testing performance across environments draws no consistent conclusions. Walker2dMassVel and Walker2dVel show minor differences between methods (see Fig. 5f), while on HalfCheetahMassVel, MPTS achieves comparable returns to ERM despite slightly higher variance. HalfCheetah verifies the marginal advantages of GDRM and ERM, whereas ReacherPos favors MPTS and DRM with reduced variance. Importantly, MPTS, serving as a DRM proxy, essentially maintains average performance comparable with ERM.
@@
In total, MPTS enjoys significant adaptation robustness benefits across most scenarios with mediate average performance, but it reserves more computational and sample efficiency than DRM. Unlike few-shot classification cases, Meta-RL mostly trades off the worst-case and average performance, similarly observed in work 33.
@@
3.3 MPTS retains multi-faced advantages beyond robustness in zero-shot continuous control